20.00 points (al Suppose tha the representatives clalm is true, and suppose that a sampe of 4 dish ownars ls randomly salactad Assuming Indapendence, use an appropnate formula to compute. (Do not round your Intermediate c calculation and ro und your answers to 4 declmal places.) Protablity 2 The probability tha more than two dish owners in the sample subecribe to Probablity 6561 at least one premium movie channel. 0037 b Siip nstt that t e representan rs c arm ts true nnd s ว nse tt at雑sample n deh Aners Is rantom sent ent 1. Nine or fewer dsh ownens in the samole subscibe to movie channe? Probahlity 2. Morc than 11 dsh owners in the sampic subocnbe to ar east one prcmium movic channc? Probablily ssuming ıntrsen ence, what t e! nbatulty that Do not round your itrmediate calculation and ound your answars to 4 dec mal laces. ar least one premium Probebility ( Suppose tha, whan we survay 20 randcmiy selected dish owners, we ind that 4 of the dish owners actualy subscribe to at bast ona premium move channel Using a protabllty you tound in this axerclse as the basis for your answer, do you beleve the Industry rapresentatves clalm Explain the daim is truc then the probabiry of tewer than 5is very large Activate Windows Book&Resources O Type here to search
Answer #1

Solution:-

b)

p = 0.10, n = 20

1) The probability that nine or fewer dish owners in the sample subscribe to atleast one premium movie channel is 0.9999.

x = 9

By applying binomial distribution:-

P(x,n) = nCx*px*(1-p)(n-x)

P(x < 9) = 0.9999

2) The probability that more than 11 fewer dish owners in the sample subscribe to atleast one premium movie channel is less than 0.00001.

x = 9

By applying binomial distribution:-

P(x,n) = nCx*px*(1-p)(n-x)

P(x > 11) = less 0.00001

3) The probability that fewer than five dish owners in the sample subscribe to atleast one premium movie channel is 0.9568.

x = 5

By applying binomial distribution:-

P(x,n) = nCx*px*(1-p)(n-x)

P(x < 5) = 0.9568